OpenStudy (anonymous):

How would I simplify or get the answer to (1/4)^-3. I need some clarification?

OpenStudy (amistre64):

a negative exponent means "inverse" or "reciprocal" (1/4)^-3 means (4/1)^3 = 4^3 = 16(4) = 64

OpenStudy (amistre64):

multiplying like bases adds eponents; dividing like bases subtracts exponents. b^3 * b^7 = b^(3+7) = b^10 b^4 / b^2 = b^(4-2) b^(4-2) = b^(4 + (-2)) b^(4+ (-2)) = b^4 * b^-2 b^4 * b^-2 = b^4 * 1/b^2